If the currency drain ratio in china and the united states is 10% of deposits, compare the money multipliers in the two countries.
1
Expert's answer
2016-04-28T09:18:02-0400
Currency drain ratio CDR = currency / deposits, money multiplier mm = 1/rr = (1 + CDR)/(rr + CDR), rr - required reserves ratio.
If the currency drain ratio in china and the united states is 10% of deposits, the money multipliers in the two countries will be mm = (1 + 0.1)/(rr + 0.1) = 1.1/(rr + 0.1), so we need to know required reserves of both countries to compare their money multipliers.